### v1

# Holographic spectral function of fermion in instantonic plasma

Abstract
Using the gauge-gravity duality, we investigate the fermionic correlation function in the D(-1)-D3 brane system which describes the instantonic plasma in holography. In this system, the charge of the D(-1) brane as the D-instanton gives the gluon condensate. To simplify the holographic setup, we first reduce briefly the ten-dimensional supergravity background produced by D(-1)-D3-branes to an equivalently five-dimensional background. Then using the standard method for computing the Green function in the AdS/CFT dictionary, we derive the equations for the fermionic correlation functions and solve them numerically with the infalling boundary condition. Our numerical results illustrate that the fermionic correlation function as spectral function includes two branches of the dispersion curves whose behavior is very close to the results obtained from the method of hard thermal loop. And the effective mass generated by the medium effect of fermion splits into two values due to the spin-dependent interactions induced by instantons. Therefore, this work in holography demonstrates the instantonic configuration is very influential to the fermionic plasmino in plasma.
Si-wen Li [^1] Email: siwenli@dlmu.edu.cn, Yi-peng Zhang [^2] Email: ypmahler111@dlmu.edu.cn, Hao-qian Li [^3] Email: lihaoqian@dlmu.edu.cn,
Department of Physics, School of Science,
Dalian Maritime University,
Dalian 116026, China
1 Introduction
In quantum chromodynamics (QCD), instanton is known as the nontrivially topological excitation of the vacuum which contributes to the thermodynamics of QCD and relates to the breaking of chiral symmetry [1, 2, 3]. Instanton is also known to consist of constituents which is called BPS (Bogomol’nyi-Prasad-Sommerfield) monopoles or dyons [4], so there have been a lot of continuous researches to connect confinement and the breaking of chiral symmetry to the instanton constituents [5, 6, 7, 8, 9, 10, 11, 12, 13]. In particular, the dynamics of fermion involving instanton e.g. the spin-dependent interaction, fermionic quantum tunneling of the instantonic vacuum, is very interesting and widely discussed in the textbooks of QCD [1, 2, 3].
However, due to the property of asymptotic freedom, it is very challenging to investigate QCD in low-energy region by using the perturbative method in quantum field theory (QFT) since QCD is strongly coupled in low-energy region. Fortunately, the gauge-gravity duality and AdS/CFT correspondence provide an alternative way to study the strongly coupled QFT [14, 15, 16] in holography. In the top-down approach, instanton is constructed simply as a geometric background with instantonic D-branes (D-instanton) [17, 18, 19, 20, 21], for example, the black D(-1)-D3 brane system can describe holographically the plasma with instanton in which the D(-1)-brane plays the role of instanton, and its dual field theory is $\mathcal{N}=4$ super Yang-Mills theory with instanton or theta angle. So there have been many works based on the holographic instanton e.g. thermodynamics of Yang-Mills theory [22], hadronic spectrum [23, 24], real-time dynamics [25], Chern-Simons theory [26, 27], chiral symmetry breaking [28], quark potential [29, 30, 31], Schwinger effect [32, 33, 34], the hadronic interaction [35, 36]. Despite all this, the study of fermion with the instanton remains to be blank to the existing works in holography although it attracts great interests as it is reviewed in [1, 2, 3] in QFT. Therefore, the goal of this work is to explore the fermionic correlation function with instanton as a start point for the further continuous investigations, since it can be evaluated by using the most fundamental principle in holography.
To our goal, the setup in this work starts with the D(-1)-D3 brane system since it describes simply the instantonic plasma in holography, then we reduce briefly the 10d IIB supergravity background produced by black D3-branes with D(-1)-branes [17, 25, 28] to an effectively five-dimensional (5d) geometric background. Afterwards, we probe a bulk fermion in the 5d geometric background and derive its covariant Dirac equation. In order to evaluate the fermionic correlation function in the dual theory, we employ the standard technique in the AdS/CFT dictionary so that obtain the differential equations for the fermionic correlators in the bulk. By imposing the infalling boundary condition at the horizon, we afterwards solve the fermionic correlation functions numerically. Our numerical calculation illustrates that, without the D-instanton, there is only one branch of the holographic dispersion curves for fermion lying on the light cone which is similar as the existing works about holographic fermions [35, 36, 37, 38, 39, 40]. In the presence of the D-instanton, the fermionic correlation function has a correction which in leading order splits the dispersion curves into two branches at small momentum, and this behavior is very close to the dispersion curves from the approach of hard thermal loop (HTL) in QFT. Furthermore, the two branches of the dispersion curves reveals that the effective bound mass of fermion in plasma has two values in the presence of the D-instanton. Remarkably it agrees qualitatively with effective potential for the spin-dependent interactions induced by instantons computed by the process of one-gluon exchange in QCD [1, 2]. Therefore, this work, as a start point, is a holographic reproduction of the properties of fermion and instanton and we believe the instantonic configuration in plasma is very influential to the fermionic plasmino.
The outline of this manuscript is given as follows. In Section 2, we collect the essentials of the 10d supergravity background produced by D3-branes with D(-1)-branes as D-instantons, and reduce it briefly to an effectively 5d gravity background. In Section 3, we demonstrate the principle in AdS/CFT dictionary to compute the holographic correlation function for spinor. Then we derive the covariant Dirac equation and the holographic counter term for spinor in order to obtain the differential equations for the fermionic correlators. In Section 4, we solve numerically the holographic fermionic correlation functions order by order with physical analysis and interpretation. The final section is the summary.
Notation
In this manuscript, the capital letters $L,M,N...$ refer to the indices of the bulk coordinate as $x^{M}$, the lowercase letters $a,b,c$ refer to the corresponding indices of the coordinate in tangent space. The gamma matrices $\gamma^{a},\Gamma^{M}$ as the generators of the Clifford algebra with elfbein $e_{M}^{a}$ are related as,
$\left\{\gamma^{a},\gamma^{b}\right\}=2\eta^{ab},\left\{\Gamma^{M},\Gamma^{M}\right\}=2g^{MN},\Gamma^{M}=e_{a}^{M}\gamma^{a},g_{MN}=e_{M}^{a}\eta_{ab}e_{N}^{a},$ (1.1)
where $g_{MN}$ refers to the metric on a manifold and $\eta_{ab}$ refers to the Minkowskian metric. The spin connection $\omega_{Mab}$ is defined as,
$\omega_{Mab}=\eta_{cb}\left(e_{a}^{N}\partial_{M}e_{N}^{c}-e_{a}^{N}\Gamma_{MN}^{K}e_{K}^{c}\right),$ (1.2)
where the affine connection $\Gamma_{MN}^{K}$ is given by,
$\Gamma_{MN}^{K}=g^{KL}\left(\partial_{M}g_{LN}+\partial_{N}g_{ML}-\partial_{L}g_{MN}\right).$ (1.3)
The Greek letters as $\mu,\nu...$refer to the indices of coordinate as $x^{\mu}$ at the holographic boundary. The the lowercase letters $i,j,k,l$ run over the spatial directions of the holographic boundary i.e. $i,j,k,l=1,2,3$. The gamma matrices presented in this work are chosen as $\gamma^{a}=\left\{\gamma^{\mu},\gamma\right\}$ as,
$\gamma^{\mu}=i\left(\begin{array}[]{cc}0&\sigma^{\mu}\\ \bar{\sigma}^{\mu}&0\end{array}\right),\gamma=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right),$ (1.4)
with $\sigma^{\mu}=\left(1,-\sigma^{i}\right),\bar{\sigma}^{\mu}=\left(1,\sigma^{i}\right),i=1,2,3$, where $\sigma^{i}$ s are the Pauli matrices.
2 The holographic background for instantonic plasma
In this section, let us briefly review the holographic background for instantonic plasma which corresponds to the system consisted of $N_{c}$ D3-branes with $N_{\mathrm{D}}$ D-instantons i.e. the D(-1)-branes in the large-$N_{c}$ limit in the type IIB string theory [17, 18, 25]. Then, we integrate out the $S^{5}$ part in order to obtain a five-dimensional (5d) background geometry for simplification.
2.1 The D3-brane background with D-instanton
The supergravity solution of $N_{c}$ D3-branes with $N_{\mathrm{D}}$ D-instantons is in general described by a 10d deformed D3-brane solution with a non-trivial Ramond-Ramond (R-R) zero form $C_{0}$. It is recognized as a marginal “bound state” of D3-branes with smeared $N_{\mathrm{D}}$ D(-1)-branes. In our notation, $N_{c}$ D3-branes are identified as color branes, hence its number $N_{c}$ denotes the color number. In the large $N_{c}$ limit, the 10d bosonic type IIB supergravity action describes the low-energy dynamics of this system which is given in string frame as,
$S_{\mathrm{IIB}}=\frac{1}{2\kappa_{10}^{2}}\int d^{10}x\sqrt{-g}\left[e^{-2\Phi}\left(\mathcal{R}+4\partial\Phi\cdot\partial\Phi\right)-\frac{1}{2}\left|F_{1}\right|^{2}-\frac{1}{2}\left|F_{5}\right|^{2}\right],$ (2.1)
where $2\kappa_{10}^{2}=\left(2\pi\right)^{7}l_{s}^{8}$ is the 10d gravity coupling constant, $l_{s},g_{s}$ is respectively the string length and the string coupling constant. We use $\Phi$ to denote the dilaton field and $F_{1,5}$ is the field strength of the R-R zero and four form $C_{0,4}$ respectively. The solution of $N_{c}$ D3-branes with $N_{\mathrm{D}}$ D-instantons is the near-horizon solution of non-extremal D3-branes with a non-trivial $C_{0}$, which in string frame reads [17, 18, 25],
$ds^{2}$ $=e^{\frac{\phi}{2}}\left\{\frac{r^{2}}{R^{2}}\left[-f\left(r\right)dt^{2}+d\mathbf{x}\cdot d\mathbf{x}\right]+\frac{1}{f\left(r\right)}\frac{R^{2}}{r^{2}}dr^{2}+R^{2}d\Omega_{5}^{2}\right\},$
$e^{\phi}$ $=1+\frac{Q}{r_{H}^{4}}\ln\frac{1}{f\left(r\right)},\ f\left(r\right)=1-\frac{r_{H}^{4}}{r^{4}},\ F_{5}=dC_{4}=g_{s}^{-1}\mathcal{Q}_{3}\epsilon_{5},$
$F_{1}$ $=dC_{0},\ C_{0}=-ie^{-\phi}+i\mathcal{C},\ \phi=\Phi-\Phi_{0},\ e^{\Phi_{0}}=g_{s},$ (2.2)
where $\mathcal{C}$ is a boundary constant for $C_{0}$, $\epsilon_{5}$ is the volume element of a unit $S^{5}$. And the associated parameters are given as,
$R^{4}=4\pi g_{s}N_{c}l_{s}^{4},\ \mathcal{Q}_{3}=4R^{4},\ Q=\frac{N_{\mathrm{D}}}{N_{c}}\frac{\left(2\pi\right)^{4}\alpha^{\prime 2}}{V_{4}}\mathcal{Q}_{3}.$ (2.3)
Here, $x^{\mu}=\left\{t,\mathbf{x}\right\}=\left\{t,x^{i}\right\},\ i=1,2,3$ refers to the 4d spacetime where the D3-branes are extended along. And $r$ is the holographic direction perpendicular to the D3-branes. The solution (2.2) is asymptotic $\mathrm{AdS}_{5}\times S^{5}$ at the holographic boundary $r\rightarrow\infty$ which describes geometrically that the D-instanton charge $N_{\mathrm{D}}$ is smeared over the worldvolume $V_{4}$ of the $N_{c}$ coincident black D3-branes homogeneously with a horizon at $r=r_{H}$. And backreaction of the D-instantons has been taken into account in the background, thus it implies $N_{\mathrm{D}}/N_{c}$ must be fixed in the large-$N_{c}$ limit. The dual theory of this background is conjectured as the 4d $\mathcal{N}=4$ super Yang-Mills theory (SYM) in a self-dual gauge field (instantonic) background or with a dynamical axion at finite temperature characterized by the parameter $Q$ [17]. Note that, in the language of hadron physics, $C_{0}$ is recognized as axion and the gluon condensate in this system can be evaluated as,
$\left\langle\mathrm{Tr}F_{\mu\nu}F^{\mu\nu}\right\rangle\simeq\frac{N_{\mathrm{D}}}{16\pi^{2}V_{4}}=\frac{Q}{\left(2\pi\alpha^{\prime}\right)^{2}R^{4}}\frac{N_{c}}{\left(2\pi\right)^{4}},\ \mu,\nu=0,1...3$ (2.4)
2.2 The effective 5d background
The $S^{5}$ part in the gravity solution (2.1) and (2.2) can be integrated out so that the background (2.2) can reduce to an 5d background equivalently. Since all the fields presented in (2.1) and (2.2) are expected to depend on the coordinates $x^{\mu},r$ only, it is possible to integrate out the $S^{5}$ part in action with
$ds_{\left(10\mathrm{d}\right)}^{2}=ds_{\left(5\mathrm{d}\right)}^{2}+ds_{S^{5}}^{2}.$ (2.5)
So the action (2.1) becomes,
$S_{\mathrm{IIB}}^{\left(5\mathrm{d}\right)}=\frac{1}{2\kappa_{5}^{2}}\int d^{5}x\sqrt{-g^{\left(5\mathrm{d}\right)}}\left[\mathcal{R}^{5\mathrm{d}}-\frac{1}{2}\partial\Phi\cdot\partial\Phi-\frac{1}{2}\left|dC_{0}\right|^{2}-2\Lambda\right],$ (2.6)
where the cosmological constant $\Lambda$ is obtained by
[…]
Accordingly, the 5d solution for the metric obtained from (2.2) satisfying the equations of motion by varying action (2.6) is
$ds_{\left(5\mathrm{d}\right)}^{2}=e^{\frac{\phi}{2}}\left\{\frac{r^{2}}{R^{2}}\left[-f\left(r\right)dt^{2}+d\mathbf{x}\cdot d\mathbf{x}\right]+\frac{1}{f\left(r\right)}\frac{R^{2}}{r^{2}}dr^{2}\right\},$ (2.8)
and the solution for $\Phi$ and $C_{0}$ remain respectively to be the corresponding formulas given in (2.2). Note that the action (2.6) describes nothing but the gravity-dilaton-axion system, and we will use the 5d effective background (2.8) in $z$ coordinate as,
$ds_{\left(5\mathrm{d}\right)}^{2}=$ $e^{\frac{\phi}{2}}\frac{R^{2}}{z^{2}}\left\{-f\left(z\right)dt^{2}+d\mathbf{x}\cdot d\mathbf{x}+\frac{dz^{2}}{f\left(z\right)}\right\},$
$f\left(z\right)=$ $1-\frac{z^{4}}{z_{H}^{4}},e^{\phi}=1-q\ln f\left(z\right),$
$z=$ $\frac{R^{2}}{r},z_{H}=\frac{R^{2}}{r_{H}},q=\frac{z_{H}^{4}Q}{R^{8}}.$ (2.9)
in which the holographic boundary is located at $z\rightarrow 0$ to continue our discussion.
3 The fermionic correlation function
In this section, we consider a probe fermion in the bulk geometry and evaluate its two-point Green function in holography by employing the standard method in the AdS/CFT dictionary. In particular, we will focus on the influence of the instanton charge denoted by $q$ in the fermionic correlation function.
3.1 General setup for the fermionic correlation function
Let us start with the principle of the AdS/CFT which illustrates that the generating functional of the dual conformal field theory (CFT) $Z_{\mathrm{CFT}}$ is equal to its gravitational partition function $Z_{\mathrm{gravity}}$ in the bulk geometry. Namely, for a spinor field $\psi$ in the 5d bulk, we have
$Z_{\mathrm{CFT}}\left[\bar{\psi}_{0},\psi_{0}\right]=Z_{\mathrm{gravity}}\left[\bar{\psi},\psi\right]\big{|}_{\bar{\psi},\psi\rightarrow\bar{\psi}_{0},\psi_{0}},$ (3.1)
with
$Z_{\mathrm{CFT}}\left[\bar{\psi_{0}},\psi_{0}\right]$ $=\left\langle\exp\left\{\int_{\partial\mathcal{M}}\left(\bar{\eta}\psi_{0}+\bar{\psi}_{0}\eta\right)d^{D}x\right\}\right\rangle,$
$Z_{\mathrm{gravity}}\left[\bar{\psi},\psi\right]$ $=e^{-S_{\mathrm{gravity}}^{ren}},$
$S_{\mathrm{gravity}}^{ren}$ $=\int_{\mathcal{M}}\mathcal{L}_{\mathrm{gravity}}^{ren}\left[\bar{\psi},\psi\right]\sqrt{-g}d^{5}x$ (3.2)
Note that, we use $\mathcal{M}$ to denote the bulk space and its holographic boundary is given at =$\partial\mathcal{M}=\left\{z\rightarrow 0\right\}$. $\psi_{0}$ refers to the boundary value of $\psi$ as a source of the boundary fermionic operator $\eta$, $\mathcal{L}_{\mathrm{gravity}}^{ren}$ denotes the renormalized Lagrangian of the bulk field $\psi$. Then follow the standard steps in quantum field theory (QFT), the one-point function i.e. average value of $\bar{\eta}$ is given as
$\left\langle\bar{\eta}\right\rangle=\frac{1}{Z_{\mathrm{CFT}}}\frac{\delta Z_{\mathrm{CFT}}}{\delta\psi_{0}}=\frac{1}{Z_{\mathrm{gravity}}}\frac{\delta Z_{\mathrm{gravity}}}{\delta\psi_{0}}=-\frac{\delta S_{\mathrm{gravity}}^{ren}}{\delta\psi_{0}}\equiv\Pi_{0},$ (3.3)
where we have used the relation (3.1). Therefore, the two-point correlation function $G_{R}$ of $\eta$ is obtained as,
$\left\langle\bar{\eta}\left(\omega,\vec{k}\right)\right\rangle=G_{R}\left(\omega,\vec{k}\right)\psi_{0},$ (3.4)
i.e.
$\Pi_{0}=G_{R}\left(\omega,\vec{k}\right)\psi_{0},$ (3.5)
where $\omega,\vec{k}$ refers to the frequency and 3-momentum of the associated Fourier modes. Altogether, it is possible to evaluate $\psi_{0},\Pi_{0}$ in order to evaluate the two-point correlation function $G_{R}$ of $\eta$ by using the classical gravity action $S_{\mathrm{gravity}}^{ren}$ in holography.
3.2 Covariant Dirac equation for a probe fermion
Keep the formulas in Section 3.1 in hand, let us investigate a probe fermion in the 5d holographic background (2.9). In general, the 5d homogeneous background metric can be written as,
$ds_{\left(5\mathrm{d}\right)}^{2}=g_{tt}dt^{2}+g_{xx}d\mathbf{x}\cdot d\mathbf{x}+g_{zz}dz^{2},$ (3.6)
where $g_{tt},g_{xx},g_{zz}$ depend on $z$ only. For a probe fermion in bulk, its dynamic is described by the Dirac action given as,
$S_{\mathrm{bulk}}=i\int d^{5}x\sqrt{-g}\bar{\psi}\left(\Gamma^{M}\nabla_{M}-m\right)\psi,$ (3.7)
so the Dirac equation reads
$\left(\Gamma^{M}\nabla_{M}-m\right)\psi=0,$ (3.8)
where the covariant derivative operator is given with the spin connection as,
$\nabla_{M}=\partial_{M}+\frac{1}{4}\omega_{Mab}\gamma^{ab},\gamma^{ab}=\frac{1}{2}\left[\gamma^{a},\gamma^{b}\right].$ (3.9)
Imposing the homogeneous metric (3.6), the Dirac operator is computed as,
$\Gamma^{M}\nabla_{M}=\frac{1}{\sqrt{-g_{tt}}}\gamma^{0}\partial_{0}+\frac{1}{\sqrt{g_{xx}}}\gamma^{i}\partial_{i}+\frac{1}{\sqrt{g_{zz}}}\gamma\partial_{z}+\frac{1}{4g_{xx}g_{tt}\sqrt{g_{zz}}}\left[3g_{tt}g_{xx}^{\prime}+g_{xx}g_{tt}^{\prime}\right]\gamma.$ (3.10)
To find a solution for the Dirac equation, we impose the ansatz for spinor $\psi$ in Dirac representation as,
$\psi$ $=\left(\begin{array}[]{c}\psi_{R}\\ \psi_{L}\end{array}\right)=\left(-gg^{zz}\right)^{-1/4}\int\frac{d^{4}p}{\left(2\pi\right)^{4}}e^{ik\cdot x}\chi\left(z,k\right),$ (3.13)
$\chi\left(z,k\right)$ $=\left[\begin{array}[]{c}\chi_{R}\left(z,k\right)\\ \chi_{L}\left(z,k\right)\end{array}\right],\chi_{R,L}=\left(\begin{array}[]{c}\chi_{R,L}^{\left(1\right)}\\ \chi_{R,L}^{\left(2\right)}\end{array}\right),k_{\mu}=\left(k_{0},k_{i}\right),$ (3.18)
into the Dirac equation (3.8), so it is simplified as,
$\left[\sqrt{\frac{g_{xx}}{g_{zz}}}\left(\gamma\partial_{z}-m\sqrt{g_{zz}}\right)+iK_{\mu}\gamma^{\mu}\right]\chi\left(z,k\right)=0,$ (3.19)
where
$K_{\mu}=\left(\sqrt{-\frac{g_{xx}}{g_{tt}}}k_{0},k_{i}\right).$ (3.20)
3.3 Construction for the correlation function
When the background metric (3.6) is inserted into (3.19), the Dirac equation reduces to two decoupled differential equations for $\chi_{R,L}$ as,
$\left(\partial_{z}+m\sqrt{g_{zz}}\right)\left[\left(K\cdot\sigma\right)^{-1}\sqrt{\frac{g_{xx}}{g_{zz}}}\left(\partial_{z}-m\sqrt{g_{zz}}\right)\chi_{R}\right]$ $=-\sqrt{\frac{g_{zz}}{g_{xx}}}\left(K\cdot\bar{\sigma}\right)\chi_{R},$
$\left(\partial_{z}-m\sqrt{g_{zz}}\right)\left[\left(K\cdot\bar{\sigma}\right)^{-1}\sqrt{\frac{g_{xx}}{g_{zz}}}\left(\partial_{z}+m\sqrt{g_{zz}}\right)\chi_{L}\right]$ $=-\sqrt{\frac{g_{zz}}{g_{xx}}}\left(K\cdot\sigma\right)\chi_{L},$ (3.21)
which can be solved analytically near the boundary $z\rightarrow 0$ as,
$\chi_{R}$ $=Cz^{1-mR}+Dz^{mR},$
$\chi_{L}$ $=Az^{-mR}+Bz^{1+mR}.$ (3.22)
where $A,B,C,D$ are constant Weyl spinor dependent on 4-momentum $k$ and the charge of the D-instanton $q$. Hence for $mR>0$, the boundary value $\chi_{0}$ of $\chi$ is given by $A$ which is the most divergent term, defined as
$\chi_{0}=\lim_{z\rightarrow 0}z^{mR}\chi=\left(\begin{array}[]{c}0\\ A\end{array}\right).$ (3.23)
Then recall the Dirac action (3.7) with (3.18), it leads to
[…]
Note that the last term in (3.24) vanishes since it is nothing but the conjugate equation of (3.19). Therefore, we can obtain a boundary action $S_{\mathrm{bdry}}$ from (3.24) as,
$S_{\mathrm{bdry}}$ $=i\int d^{4}x\left(\bar{\chi}\gamma\chi\right)|_{z\rightarrow 0}$
$=i\int d^{4}x\left(D^{\dagger}A+C^{\dagger}Az^{1-2m}\right)|_{z\rightarrow 0},$ (3.25)
which implies the holographic counterterm $S_{\mathrm{CT}}$ at boundary is,
$S_{\mathrm{CT}}=-i\int d^{4}xC^{\dagger}Az^{1-2m}|_{z\rightarrow 0},$ (3.26)
giving the holographic renormalized action as,
$S_{\mathrm{gravity}}^{ren}$ $=S_{\mathrm{bdry}}+S_{\mathrm{CT}}$
$=i\int d^{4}xD^{\dagger}A.$ (3.27)
Picking up (3.27) with (3.3) and (3.5), the correlation function satisfies
$D=-G_{R}A,$ (3.28)
in the Weyl representation. This is not surprised since the boundary value $\psi_{0}$ of $\psi$ is in fact a chiral spinor. On the other hand, the correlation function in the $2\times 2$ matrix must be able to be written by the combination of the Pauli matrices as,
$G_{R}=\alpha+\beta k_{i}\sigma^{i},$ (3.29)
we can set that the momentum is along $x^{3}$ direction as $k_{\mu}=\left(-\omega,0,0,\mathrm{k}\right)$ due to the rotation symmetry in $\mathbb{R}^{3}=\left\{x^{i}\right\}$, so that the correlation function $G_{R}$ is expected to be diagonal, as
$G_{R}=\left(\begin{array}[]{cc}G_{R}^{1,1}&0\\ 0&G_{R}^{2,2}\end{array}\right).$ (3.30)
Figure 1: The real and imaginary parts of zero-th order correlation functions $G_{R,\alpha}^{\left(0\right)}$ as functions of $\omega,\mathrm{k}$. The parameter is chosen as $q=0,m=0.01,z_{H}=1.$
Figure 2: Density plots of the zero-th order correlation functions $G_{R,\alpha}^{\left(0\right)}$ with the parameters as $q=0,m=0.01,z_{H}=1.$ The white regions refer to the peaks in the correlation functions.
To finalize this section, we introduce the ratio defined as
$G_{\alpha}=\left(-1\right)^{\alpha+1}\frac{\chi_{R}^{\left(\alpha\right)}}{\chi_{L}^{\left(\alpha\right)}},\ \alpha=1,2.$ (3.31)
According to (3.23) - (3.28), the correlation function is given by
$G_{R}^{\alpha,\alpha}=\left(-1\right)^{\alpha}\lim_{z\rightarrow 0}z^{-2m}G_{\alpha},$ (3.32)
and the equations for the ratios can be obtained by recombining the components of the spinor $\chi$ with respect to the Dirac equation (3.19). As a result, one can find the equations for the ratios are given as,
$\sqrt{\frac{g_{xx}}{g_{zz}}}G_{\alpha}^{\prime}=\left(-1\right)^{\alpha}u-\mathrm{k}+\left[\mathrm{k}+\left(-1\right)^{\alpha}u\right]G_{\alpha}^{2}+2m\sqrt{g_{xx}}G_{\alpha},$ (3.33)
where “” refers to the derivative with respect to $z$, and
$u=\omega\sqrt{-\frac{g_{xx}}{g_{tt}}}=\frac{\omega}{\sqrt{f}}.$ (3.34)
Therefore the correlation function could be evaluated numerically by solving (3.33) with the infalling boundary condition $G_{\alpha}|_{z=z_{H}}=\left(-1\right)^{\alpha}i$ obtained by analyzing its asymptotic behavior near $z=z_{H}$ as [35, 36, 37, 38, 39]. Furthermore, when the D-instanton is taken into account, the correlation function given in (3.32) should also depend on the charge density of the D-instanton denoted as $q$. In order to evaluate exactly the $q$-dependence in the correlation function, we consider its leading order correction by expand $G_{\alpha}$ as,
$G_{\alpha}=\mathcal{G}_{\alpha}+q\xi_{\alpha},$ (3.35)
where $\mathcal{G}_{\alpha}$ is the zero-th order solution satisfying the equations presented in (3.33) with $q=0$ as (there is no sum for the repetitive indices),
$\sqrt{f}\mathcal{G}_{\alpha}^{\prime}=\left(-1\right)^{\alpha}u-\mathrm{k}+\left[\mathrm{k}+\left(-1\right)^{\alpha}u\right]\mathcal{G}_{\alpha}^{2}+\frac{2mR}{z}\mathcal{G}_{\alpha}.$ (3.36)
And $\xi_{\alpha}$ is the leading order correction satisfying the equations
$\sqrt{f}\xi_{\alpha}^{\prime}=$ $\left[\mathrm{k}+\left(-1\right)^{\alpha}u\right]2\mathcal{G}_{\alpha}\xi_{\alpha}+\frac{2mR}{z}\left[\xi_{\alpha}-\frac{1}{4}\mathcal{G}_{\alpha}\ln f\right].$ (3.37)
By analyzing the asymptotic behavior of (3.36), we find the infalling boundary condition for $\mathcal{G}_{\alpha}$ could be $\mathcal{G}_{\alpha}|_{z=z_{H}}=\left(-1\right)^{\alpha}i$ which is exactly same as the infalling boundary condition for $G_{\alpha}$. Therefore, the boundary condition for $\xi_{\alpha}$ is expected as $\xi_{\alpha}|_{z=z_{H}}=0$. Altogether, we will solve (3.36) and (3.37) numerically in order to evaluate the associated correlation function $G_{R}$ as the fermionic spectral function in holography.
4 The numerical analysis
Figure 3: The real and imaginary parts of leading order correlation functions $G_{R,\alpha}^{\left(1\right)}$ as functions of $\omega,\mathrm{k}$ with $m=0.01,z_{H}=1.$
Figure 4: Density plots of leading order correlation functions $G_{R,\alpha}^{\left(1\right)}$ as functions of $\omega,\mathrm{k}$ with $m=0.01,z_{H}=1.$ The white regions refer to the peaks in the correlation functions
Figure 5: The real and imaginary parts of leading order correlation functions $G_{R,\alpha}^{\left(1\right)}$ as functions of $\omega$ with various $\mathrm{k}$ and $m=0.01,z_{H}=1$. There are always two peaks for various $\mathrm{k}$ in the correlation function which correspond to the two branches of the dispersion curves.
We write down the correlation functions presented in (3.32) as the combinations of its zero-th order $G_{R,\alpha}^{\left(0\right)}$ and first order solution $G_{R,\alpha}^{\left(1\right)}$ as,
$G_{R}^{\alpha,\alpha}$ $=G_{R,\alpha}^{\left(0\right)}+qG_{R,\alpha}^{\left(1\right)},$
$G_{R,\alpha}^{\left(0\right)}$ $=\left(-1\right)^{\alpha}\lim_{z\rightarrow 0}z^{-2m}\mathcal{G}_{\alpha},$
$G_{R,\alpha}^{\left(1\right)}$ $=\left(-1\right)^{\alpha}\lim_{z\rightarrow 0}z^{-2m}\xi_{\alpha},$ (4.1)
where $\mathcal{G}_{\alpha},\xi_{\alpha}$ are coefficients independent on $q$ and they can be solved numerically through the equations (3.36) and (3.37). The corresponding numerical results for $\mathcal{G}_{\alpha},\xi_{\alpha}$ are given in Figure 1 - Figure 5.
Let us compare our numerical results with the calculations by using the method of hard thermal loop (HTL) in thermal QCD. The approach of HTL with vanished chemical potential illustrates the dispersion curves of fermions (plasmino) is  [^4] We pick up the modes of positive frequency with $\omega>0$ and momentum $\mathrm{k}>0$, the case for the negative frequency can be obtained by replacing $\omega\rightarrow-\omega$.[41],
$\omega\left(\mathrm{k}\right)$ $\simeq m_{f}\pm\frac{1}{3}\mathrm{k}+\frac{1}{3m_{f}}\mathrm{k}^{2},\ \mathrm{k}\ll 1;$
$\omega\left(\mathrm{k}\right)$ $\simeq\mathrm{k},\ \mathrm{k}\gg 1,$ (4.2)
where $m_{f}$ is the effective mass generated by the medium effect of fermion as,
$m_{f}=\sqrt{\frac{C_{F}}{8}}g_{\mathrm{YM}}T.$ (4.3)
Here $g_{\mathrm{YM}}$ refers to the Yang-Mills coupling constant and $C_{F}$ is a constant suggested to be $C_{F}=4/3$ for fundamental quarks or $C_{F}=1$ for electron. Accordingly, we can see the dispersion curves presented in zero-th order correlation function given in Figure 1 and Figure 2 cover basically the behavior of dispersion curves from HTL at large momentum. For the small momentum, while the zero-th order dispersion curves may be close to the relation given in (5), it does not illustrate two branches of the dispersion curves as some related works in holography [35, 36, 37, 38, 39]. The reason may be that, according to the gravity/fluid correspondence [42, 43], the plasma described holographically by metric (2.8) with $q=0$ corresponds to conformally ideal fluid of $\mathcal{N}=4$ super Yang-Mills theory which is different from QED or QCD plasma. In particular, the interaction of the plasmino or the dissipation of the medium is not taken into account in the ideal fluid. While this issue may be expected to be figured out by taking into account the dissipation of the fluid somehow in the holographic background (2.8), we will focus on the influence on the instanton in this work.
For the case $q>0$, the correlation function has a correction term $G_{R,\alpha}^{\left(1\right)}$ given by $\xi_{\alpha}$ as it is illustrated in Figure 3 - Figure 5. For $\omega>0$ and $\mathrm{k}>0$, it is clear to see that the dispersion curves have two branches at small momentum and they trend to merge for large momentum. By taking a look at the value of the correction terms, we can find the correction $G_{R,\alpha}^{\left(1\right)}$ is large/small for small/large momentum respectively which has an opposite behavior to the zero-th correlation function $G_{R,\alpha}^{\left(0\right)}$. Altogether, for $q>0$, the first order correction $G_{R,\alpha}^{\left(1\right)}$ dominates the behavior of the dispersion curves at small momentum and it always illustrates two branches. At large momentum, zero-th order function $G_{R,\alpha}^{\left(0\right)}$ dominates the behavior of the dispersion curves which displays only one branch of the dispersion curves on the light cone. Therefore, the dispersion curves given by the total spectral function $G_{R}^{\alpha,\alpha}$ are very close to the relation for HTL (4.2).
Besides, our numerical results also illustrate that, i.e. for $\omega,\mathrm{k}>0$, the effective mass $m_{f}=\omega\left(\mathrm{k}=0\right)$ splits into two values as it is displayed in Figure 4. This may relate remarkably to the properties of the fermionic spin, since it is known that there is spin-dependent interactions among fermions induced by instantons in QCD [1, 2]. According to the calculations by one-gluon exchange, the massive quark potential dependent on the quantum number of spin in the presence of instanton, so the total mass of the plasmino including quark interaction splits into two values which leads to two branches in the dispersion. In this sense, our numerical evaluation might be a holographic reproduction of instanton-induced interaction with spin. However, for the massless case, the spin-dependent interaction induced by instantons trend to be vanished, so it could be why the spectral function of massless fermion is not affected by the instanton charge. On the other hand, some existing works about D-instanton in gauge-gravity duality [24, 28, 29, 30, 31, 32, 33, 35, 36] also support that the property of mass for a massless hadron in plasma is not affected by the instanton, which could thus be a holographic confirmation.
5 Summary
In this work, we study the correlation function of fermion in the D(-1)-D3 brane system which describes holographically the hot plasma with instantonic configuration. To simplify the setup, we briefly reduce the 10d supergravity background to an 5d background of gravity-dilaton-axion system. By probing a spinor in the 5d bulk, we derive the equations for the holographic correlation function from the Dirac equation, then demonstrate the numerical calculations with the infalling boundary condition. In particular, we expand the correlation function as zero-th order solution plus first order solution. Our numerical results illustrate that the first order correction to the total correlation function including two branches of the fermionic dispersion curves is dominated for small momentum, while the zero-th order correlation function dominates the behavior of the fermionic dispersion curves at large momentum giving only one merged dispersion curves on light cone. This behavior is out of reach without the presence of the instanton and it is very close to the fermionic dispersion curves obtained by HTL. Furthermore, we also find the effective mass generated by the medium effect of fermion splits into two values due to the spin-dependent interactions among fermions induced by instantons which has be studied in the framework of QCD with instanton [1, 2]. Altogether, it means seemingly the instantonic configuration in the plasma is very influential for the spectral function of the plasmino.
Acknowledgements
We would like to thank Prof.Lin Shu for helpful discussion. This work is supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12005033 and the Fundamental Research Funds for the Central Universities under Grant No. 3132024192.